P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

1/18 ifmman.tex – June 24, 2001

TU Darmstadt – Institute of Mechanics – AG4

http://coulomb.mechanik.tu-darmstadt.de/

ifmslide.sty v0.44 November 13, 2000

Enhanced presentations with (PDF)L^{A}TEX combining the texpower, hyperref and seminar-packages

Thomas Emmel

emmel@mechanik.tu-darmstadt.de

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

About ifmslide.sty

Early in May 2000 ...

OK! Cancel this stirring story about my problems generating printouts for my presentation. What can ifmslide.sty do for you?

You like to do nice presentations of your business with beamers or on a screen?

You are not sure that all these things work when you need them? You feel better with a stack of printed slides in your bag?

ifmslide.sty provides both: getting a presentation with PDFL^{A}TEX and printouts with L^{A}TEX and - as a side effect - simple production of your slides using your favourite DVI-viewer.^{a} You don’t need PPower4 to get all these nice effects with page-transitions and stepwise building of the pages. All you need is texpower [1] and hyperref [2] for the links and buttons etc. ifmslide.sty makes use of the special features of the classes seminar and powersem (part of TEXPower).

2/18 ifmman.tex – June 24, 2001

aThis is not pdfslide [3] as it seems to be in a first view. It was indeed developed from that style but now completely

rewritten with many new features implemented and functionally extented. pdfslide cannot produce printouts

without generating a PDF-file and rescaling the output with all difficulties.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer).

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features. ... panel position is free: right, left, bottom, top and outside of the slide! You can

change it in the document as you like.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features. ... panel position is free: right, left, bottom, top and outside of the slide! You can

change it in the document as you like. ... position of the buttons and the logo is completey free!

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features. ... panel position is free: right, left, bottom, top and outside of the slide! You can

change it in the document as you like. ... position of the buttons and the logo is completey free! ... changing the size of your slide...

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features. ... panel position is free: right, left, bottom, top and outside of the slide! You can

change it in the document as you like. ... position of the buttons and the logo is completey free! ... changing the size of your slide... ... free choice of button-design, background for the panel and the frame...

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

3/18 ifmman.tex – June 24, 2001

Features

... produces DVI (L^{A}TEX) for printouts and PDF (PDFL^{A}TEX) for direct presentation. ... DVI-Version with extra margins for the printer. ... draft-mode with simple frames instead of colored boxes (easier to display and

debug your slides with your DVI-viewer). ... direct use of bookmarks to navigate in the PDF-version. ... local or global configuration file ifmslide.cfg for colors and many features. ... panel position is free: right, left, bottom, top and outside of the slide! You can

change it in the document as you like. ... position of the buttons and the logo is completey free! ... changing the size of your slide... ... free choice of button-design, background for the panel and the frame... ... you can use most of the features of the seminar-package: magnification etc.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Options

The following options are provided:

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Options

The following options are provided:

draft shows colored boxes as white boxes with frames, efficient for debugging and writing the slides. The pause-command is displayed as a small orange box.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Options

The following options are provided:

draft shows colored boxes as white boxes with frames, efficient for debugging and writing the slides. The pause-command is displayed as a small orange box.

display TEXPower-option: all effects are turned off if not set!

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Options

The following options are provided:

draft shows colored boxes as white boxes with frames, efficient for debugging and writing the slides. The pause-command is displayed as a small orange box.

display TEXPower-option: all effects are turned off if not set!

ams, cnav, cnavo, stmo turn the navigation buttons on and sets the style. Just try it...

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Options

The following options are provided:

draft shows colored boxes as white boxes with frames, efficient for debugging and writing the slides. The pause-command is displayed as a small orange box.

display TEXPower-option: all effects are turned off if not set!

ams, cnav, cnavo, stmo turn the navigation buttons on and sets the style. Just try it...

button use definable buttons together with the last option instead of just putting a box around the signs.

4/18 ifmman.tex – June 24, 2001

colorinfo from older versions will be ignored, contnav and amsnav are changed to cnav and ams

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Pagestyles

sidebar, panel Switch the navigation panel on. plain, myheadings, headings are redefined to it. You can put the panel at all positions with \panelposition

5/18 ifmman.tex – June 24, 2001

pagestyle title is changed to sidebar and background is deleted, because of the \background-command

Pagestyles

empty No panel – only a frame and background.

But you can put a panel outside of the frame with \panelposition which is not present in L^{A}TEX-mode and outside of the normal papersize in PDFL^{A}TEX-mode!

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Pagestyles

panelposition

\panelposition{left/right/top/bottom} changes the position of the panel to

the given value. The slide is recalculated for every position...

\panelposition{outsidebottom} set a panel outside of the paper!

\panelposition{empty} set no panel which is in practice useful with

\pagestyle{empty} only.

⪦ ⪧ ⦇⪦ ⪧⦈ ⪧ ⪧ ⪦ ⪦ ⦵ _{?} _{i} ⧈ _{P}

Panel on top

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦  _{ifmman.tex – June 24, 2001}_{9/18}

Panel left

⪦ ⪧ ⦇⪦ ⪧⦈ ⪧ ⪧ ⪦ ⪦ ⦵ _{?} _{i} ⧈ _{P}

Panel on bottom

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Panel right Panel outsidebottom (you cannot see it on the printed slides)!

ifmman.tex – June 24, 2001_{13/18}  _{⪦ ⪧ ⦇⪦ ⪧⦈ ⪧ ⪧ ⪦ ⪦ ⦵ _{?} _{i} ⧈ _{P}}

Buttons and logo

If you don’t like to put the buttons and the logo into the panel, you can release them

with \releaselogo and \releasebutton. \catchbutton catches them ;-).
\buttonsize(width_{max},height_{max})(distance_{h},distance_{v})] gives the maxi-

mal size of the buttons (aspectratio is kept) and the distance between the buttons without length. Keep in mind that the frame and the panel is placed into a pictureenvironment with \unitlength=1mm and all positions given in mm. \freebutton(pos_{x},pos_{y})[h/v] is the position of the buttons and the direction [h] for horizontal and [v] for vertical buttons. For horizontal buttons the anchor is the lower right point of the right button and for vertical buttons the point between the lowest buttons. Use \nobuttons to disable buttons all over (e.g. for slides). \freelogo(pos_{x},pos_{y})[width] the anchor for the position is the lower left point of the logo, the width should be clear. If you think that the pagecounter is too lonely in PDF-mode switch him off with

\pagecounter[off/on].

ifmman.tex – June 24, 2001_{14/18}  _{⪦ ⪧ ⦇⪦ ⪧⦈ ⪧ ⪧ ⪦ ⪦ ⦵ _{?} _{i} ⧈ _{P}}

useful application for this feature....

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Background

Older versions of ifmslide.sty provide \pagestyle{background} which makes

objectively no sense, due to the fact that one needs a background but no panel or vice versa. The \background{filename} command substitute it together with the

switches \showbackground and \hidebackground.

A new feature for the panel-background is:

\paneltile[on/off](cut_{llx},cut_{lly},cut_{urx},cut_{ury}){filename}

which is now fully experimental which means that it can change it’s definition and possibilities in further versions. Presently it works only for vertical panels. It takes a picture^{a} and fill the panel with tiles of it. It takes one tile in x-direction and so many as needed to fill in y-direction. There is still a rest which is filled by a part of the picture defined by the four cut-values... That did not work in all cases and in PDF-mode the rest of the tile is printed outside the visible area, but it is pre-alpha so what.

15/18 ifmman.tex – June 24, 2001

ajpeg etc. in PDF-mode and it’s EPS-version for DVI-mode (produced with some tool like xv)

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Useful commands

\setInsideMargin{left}{right}{top}{bottom} set the extra margins inside the

frame. This is the easiest way to improve your output.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Useful commands

\setInsideMargin{left}{right}{top}{bottom} set the extra margins inside the

frame. This is the easiest way to improve your output. Take \headskip=length to set the \section-title to the correct vertical position.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Useful commands

\setInsideMargin{left}{right}{top}{bottom} set the extra margins inside the

frame. This is the easiest way to improve your output.

Take \headskip=length to set the \section-title to the correct vertical position.

Make own buttons and use them with \OnButton{file}, \OffButton{file} and

\DraftButton{file}.

There is a tiny script called genbutton coming with ifmslide.sty which can be used to change the color of the standard buttons (button1..) and write new buttons.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Hints

The file ifmslide.cfg can be changed to set up global or local values such as colors, the logo, backgrounds, sizes, the baseurl of your company, buttons, margins ... without changing ifmslide.sty itself.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Hints

The file ifmslide.cfg can be changed to set up global or local values such as colors, the logo, backgrounds, sizes, the baseurl of your company, buttons, margins ... without changing ifmslide.sty itself.

Make use of \setslidelength{somelength}{somesize}, \semcm and \semin as

described in seminar.sty.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Hints

The file ifmslide.cfg can be changed to set up global or local values such as colors, the logo, backgrounds, sizes, the baseurl of your company, buttons, margins ... without changing ifmslide.sty itself.

Make use of \setslidelength{somelength}{somesize}, \semcm and \semin as

described in seminar.sty.

Change font sizes with \slidesmag{mag} and \ptsize{size} from seminar.sty.

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Hints

The file ifmslide.cfg can be changed to set up global or local values such as colors, the logo, backgrounds, sizes, the baseurl of your company, buttons, margins ... without changing ifmslide.sty itself.

Make use of \setslidelength{somelength}{somesize}, \semcm and \semin as

described in seminar.sty.

Change font sizes with \slidesmag{mag} and \ptsize{size} from seminar.sty.

Produce an EPS and a PDF-version^{a} of your pictures and include them with

\includegraphics[height=5\semcm,width=4\semcm]{file}^{b}

P ⧈_{i} _{?} ⦵ ⪦ ⪦ ⪧ ⪧ ⪧⦈ ⦇⪦ ⪧ ⪦

Hints

The file ifmslide.cfg can be changed to set up global or local values such as colors, the logo, backgrounds, sizes, the baseurl of your company, buttons, margins ... without changing ifmslide.sty itself.

Make use of \setslidelength{somelength}{somesize}, \semcm and \semin as

described in seminar.sty.

Change font sizes with \slidesmag{mag} and \ptsize{size} from seminar.sty.

Produce an EPS and a PDF-version^{a} of your pictures and include them with

\includegraphics[height=5\semcm,width=4\semcm]{file}^{b}

Further information about \step, \pause, \href, slides and other stuff can be found in the manuals for the packages texpower [1] , hyperref [2] and seminar.

17/18 ifmman.tex – June 24, 2001

aepstopdf is a good tool to convert eps and the graphicx package knows many formats...

bfile is the filename without suffix

References

[1] texpower-Package: Stephan Lehmke, (Stephan.Lehmke@cs.uni-dortmund.de), University of Dortmund or CTAN in the future.

[2] hyperref-Package: Sebastian Rahtz, www.tug.org

[3] pdfslide.sty: C. V. Radhakrishnan, (cvr@river-valley.com), CTAN-archive.